# ProtoECGNet: Case-Based Interpretable Deep Learning for Multi-Label ECG Classification with Contrastive Learning

### v2

@@ Abstract
Deep learning-based electrocardiogram (ECG) classification has shown impressive performance but clinical adoption has been slowed by the lack of transparent and faithful explanations. Post hoc methods such as saliency maps may fail to reflect a model’s true decision process. Prototype-based reasoning offers a more transparent alternative by grounding decisions in similarity to learned representations of real ECG segments—enabling faithful, case-based explanations. We introduce ProtoECGNet, a prototype-based deep learning model for interpretable, multi-label ECG classification. ProtoECGNet employs a structured, multi-branch architecture that reflects clinical interpretation workflows: it integrates a 1D CNN with global prototypes for rhythm classification, a 2D CNN with time-localized prototypes for morphology-based reasoning, and a 2D CNN with global prototypes for diffuse abnormalities. Each branch is trained with a prototype loss designed for multi-label learning, combining clustering, separation, diversity, and a novel contrastive loss that encourages appropriate separation between prototypes of unrelated classes while allowing clustering for frequently co-occurring diagnoses. We evaluate ProtoECGNet on all 71 diagnostic labels from the PTB-XL dataset, demonstrating competitive performance relative to state-of-the-art black-box models while providing structured, case-based explanations. To assess prototype quality, we conduct a structured clinician review of the final model’s projected prototypes, finding that they are rated as representative and clear. ProtoECGNet shows that prototype learning can be effectively scaled to complex, multi-label time-series classification, offering a practical path toward transparent and trustworthy deep learning models for clinical decision support.
@@ 1 Introduction
Deep learning (DL) has achieved strong performance across a wide range of diagnostic and predictive tasks in medicine. Aggarwal et al. (2021); Petmezas et al. (2022); Khera et al. (2024); Oliveira et al. (2023); Barnes et al. (2023); Sethi et al. (2025); Sun et al. (2021). In cardiology, DL-based electrocardiogram (ECG) interpretation is of great interest because ECGs are central to diagnosing many diseases like arrhythmias, myocardial infarction, and structural heart disease Carrington et al. (2022); Birnbaum et al. (2014). DL models have demonstrated strong, and in some cases, cardiologist-level performance in ECG classification Hannun et al. (2019); Elias et al. (2022); He et al. (2023); Ouyang et al. (2024); Trivedi et al. (2025); Yuan et al. (2023), but their clinical deployment would be accelerated with increased transparency and trustworthiness in model predictions Goettling et al. (2024). An example of a blackbox prediction for an ECG with an anteroseptal myocardial infarction (ASMI) is shown in Figure 1A.
Post hoc explainability methods—such as saliency maps and attention-based visualizations—are commonly used to interpret black-box deep learning models Rudin (2019); Adebayo et al. (2020). However, these methods generate outputs that are not necessarily aligned with the model’s decision process. Prior studies have shown that saliency maps can be unstable, non-reproducible, and misaligned with human reasoning, particularly in medical domains Adebayo et al. (2020); Alvarez-Melis and Jaakkola (2018); Turbé et al. (2023). Most importantly, simply highlighting the parts of an ECG the model focused on is not the same as explaining why it made a specific diagnosis—as illustrated in Figure 1B. These limitations have led to growing interest in self-explaining models, where interpretability is embedded directly into the model architecture Rudin (2019); Tonekaboni et al. (2019).
@@
    A multi-branch model architecture aligned with clinical reasoning. ProtoECGNet consists of three specialized prototype branches: (1) a 1D rhythm model with global prototypes designed to capture long-range temporal patterns, (2) a 2D morphology model with time-localized prototypes that leverage inter-lead spatial structure to identify focal waveform abnormalities, and (3) a 2D global model with full-duration, full-lead prototypes for detecting diffuse or distributed ECG abnormalities. Each branch is trained independently on a disjoint subset of diagnostic labels, and their outputs are either aggregated via macro-averaging or fused using a learned classifier trained to predict all 71 labels in the PTB-XL dataset.
@@
    Empirical evaluation of prototype quality by clinicians. To assess the interpretability of the learned prototypes, we conducted a structured review in which two physicians independently rate all projected prototypes from the final model on multiple quality criteria, including clarity and class representativeness. This evaluation provides initial evidence that the model’s explanations align with clinical expectations.
@@ 2.2 Contrastive Learning
Our approach is inspired by supervised contrastive learning, which has been widely used to structure encoder representations. SupCon Khosla et al. (2021) and SimCLR Chen et al. (2020) promote similarity among positive pairs and dissimilarity among negatives using log-softmax objectives over large batches. Multi-label extensions, such as MulSupCon Zhang and Wu (2024), weight similarity based on label overlap. However, these instance-level methods are not applicable to our setting, where the goal is to organize a fixed set of prototype vectors with known class assignments—not to embed unlabeled instances.
@@ 3.1 Dataset and Label Grouping
We used the PTB-XL dataset, a publicly available collection of 21,799 10-second, 12-lead ECGs from 18,869 patients sampled at both 500 Hz and 100 Hz Wagner et al. (2020). Each ECG was annotated with one or more of 71 SCP-ECG diagnostic labels spanning arrhythmias, conduction disorders, infarction patterns, and morphological abnormalities. We retained the original dataset split from Wagner et al. (2020)—using folds 1-8 for training and fold 9 for validation. We reported performance metrics on fold 10, using it as a hold-out test set. To align with the dataset’s benchmarks, the term-centric macro AUROC from Strodthoff et al. (2021) was chosen as the primary performance metric. Appendix F details how ECGs were visualized for figures and prototype review.
The dataset contains several label groupings (e.g., diagnostic superclasses), but these were unsuitable for our desired training process. Consequently, two physicians—one board-certified in cardiology and the other in internal medicine—grouped the 71 labels into three clinically meaningful prototype categories based on the type of visual reasoning required for diagnosis: (1) rhythm-based diagnoses—require temporal pattern analysis across full-length ECG signals, often discernible from a single lead; (2) morphology-based diagnoses—require localized waveform shape or inter-lead comparisons over short time intervals; (3) global diagnoses—require full-lead patterns spanning the full ECG duration. In total, 16 diagnoses were grouped into the 1D rhythm branch, 52 into the 2D morphology branch, and 3 into the 2D global branch. These groupings are detailed in Appendix G.
@@ 3.2 Preprocessing
As the lower end of the frequency range for a normal ECG is 0.5 Hz Zheng et al. (2020), we applied a first-order Butterworth high-pass filter with a 0.5 Hz cutoff. No low-pass filter was applied, as we used the 100 Hz samples from the dataset.
@@ 3.3 Model Architecture
This branch was designed to capture long-range temporal dependencies characteristic of rhythm abnormalities. We adopted the ResNet1D-18 architecture from Strodthoff et al. (2021), operating on input ECGs of shape $(12\times 1000)$. The architecture consisted of an initial strided convolution, followed by four residual blocks with increasing channel depth. Feature maps were pooled with an adaptive average pooling layer to produce fixed-length latent representations per lead. These representations formed the prototype matching space for the 1D rhythm model.
@@
To detect localized waveform abnormalities that depend on spatial relationships across ECG leads—such as ST elevation or pathological Q waves—we used a 2D convolutional neural network that treats the 12-lead ECG as a spatial-temporal matrix. Input ECGs are shaped as $(1\times 12\times 1000)$, where the vertical dimension corresponded to the 12 leads and the horizontal dimension to time. The model used a modified ResNet18 He et al. (2015) backbone in which the first convolutional layer was adapted to accept single-channel 2D inputs with a filter size of $(12\times 7)$. Pretrained ImageNet weights were loaded for the remaining layers. To preserve temporal resolution, the global average pooling layer was removed, resulting in a final feature map of shape $(512\times 1\times 32)$, where the temporal axis was down sampled from 1000 to 32 steps. This branch used partial prototypes that spanned a small temporal window (3 units, approximately 0.94 seconds at 100 Hz) and were applied in a sliding fashion across the latent time dimension. During inference, the model computed similarity scores between each partial prototype and all possible time-localized windows of the test ECG’s latent feature map. Top-$k$ pooling was then applied to aggregate the $k$ most activated positions into a final similarity score per prototype.
@@
[…]
@@ 3.4 Prototype Loss Function
Together, our modifications to the clustering and separation loss terms extended their logic from requiring an exact label match between a prototype and a training example, to allowing prototypes to be attracted to any example that contains their assigned class—regardless of what other labels are present. This enabled clustering in the presence of multi-label supervision and implicitly supported overlap among frequently co-occurring diagnoses. However, these terms do not directly structure the relationships between prototypes themselves. To more explicitly shape the geometry of the prototype space based on diagnostic co-occurrence, we introduced a contrastive loss applied at the level of prototype-prototype similarity. This encouraged prototypes assigned to frequently co-occurring classes to remain similar in the latent space, while discouraging similarity between prototypes associated with rarely co-occurring diagnoses.
@@
Our contrastive loss was inspired by the Jaccard Similarity Contrastive Loss (JSCL) proposed by Lin et al. (2023) in the context of multi-label text classification. JSCL weights the contrastive objective between sample pairs by the Jaccard similarity of their label sets, allowing for a soft notion of positive and negative pairs in multi-label settings. We adopted this core idea—using Jaccard similarity to scale pairwise contrastive terms—and adapted it to prototype learning by computing similarity between learned prototype vectors. Specifically, we precomputed a Jaccard-based co-occurrence matrix $C\in[0,1]^{P\times P}$ over the training label set and used it to weight the pairwise similarity score between prototypes.

### v5

@@ Abstract
Deep learning-based electrocardiogram (ECG) classification has shown impressive performance but clinical adoption has been slowed by the lack of transparent and faithful explanations. Post hoc methods such as saliency maps may fail to reflect a model’s true decision process. Prototype-based reasoning offers a more transparent alternative by grounding decisions in similarity to learned representations of real ECG segments—enabling faithful, case-based explanations. We introduce ProtoECGNet, a prototype-based deep learning model for interpretable, multi-label ECG classification. ProtoECGNet employs a structured, multi-branch architecture that reflects clinical interpretation workflows: it integrates a 1D CNN with global prototypes for rhythm classification, a 2D CNN with time-localized prototypes for morphology-based reasoning, and a 2D CNN with global prototypes for diffuse abnormalities. Each branch is trained with a prototype loss designed for multi-label learning, combining clustering, separation, diversity, and a novel contrastive loss that encourages appropriate separation between prototypes of unrelated classes while allowing clustering for frequently co-occurring diagnoses. We evaluate ProtoECGNet on all 71 labels from the PTB-XL dataset, demonstrating competitive performance relative to state-of-the-art black-box models while providing structured, case-based explanations. To assess prototype quality, we conduct a structured clinician review of the final model’s projected prototypes, finding that they are rated as representative and clear. ProtoECGNet shows that prototype learning can be effectively scaled to complex, multi-label time-series classification, offering a practical path toward transparent and trustworthy deep learning models for clinical decision support.
@@ 1 Introduction
Deep learning (DL) has achieved strong performance across a wide range of diagnostic and predictive tasks in medicine Aggarwal et al. (2021); Petmezas et al. (2022); Khera et al. (2024); Oliveira et al. (2023); Barnes et al. (2023); Sethi et al. (2025); Sun et al. (2021). In cardiology, DL-based electrocardiogram (ECG) interpretation is of great interest because ECGs are central to diagnosing many diseases like arrhythmias, myocardial infarction, and structural heart disease Carrington et al. (2022); Birnbaum et al. (2014). DL models have demonstrated strong, and in some cases, cardiologist-level performance in ECG classification Hannun et al. (2019); Elias et al. (2022); He et al. (2023); Ouyang et al. (2024); Trivedi et al. (2025); Yuan et al. (2023), but their clinical deployment would be accelerated with increased transparency and trustworthiness in model predictions Goettling et al. (2024). An example of a blackbox prediction for an ECG with an anteroseptal myocardial infarction (ASMI) is shown in Figure 1A.
Post hoc explainability methods—such as saliency maps and attention-based visualizations—are commonly used to interpret black-box deep learning models Rudin (2019); Adebayo et al. (2020). However, these methods generate outputs that are not necessarily aligned with the model’s decision process. Prior studies have shown that saliency maps can be unstable, non-reproducible, and misaligned with human reasoning, particularly in medical domains Adebayo et al. (2020); Alvarez-Melis and Jaakkola (2018); Turbé et al. (2023). Most importantly, simply highlighting the parts of an ECG that the model focused on is not the same as explaining why it made a specific diagnosis—as illustrated in Figure 1B. These limitations have led to growing interest in self-explaining models, where interpretability is embedded directly into the model architecture Rudin (2019); Tonekaboni et al. (2019).
@@
    A multi-branch model architecture aligned with clinical reasoning. ProtoECGNet consists of three specialized prototype branches: (1) a 1D rhythm model with global prototypes designed to capture long-range temporal patterns, (2) a 2D morphology model with time-localized prototypes that leverage inter-lead spatial structure to identify focal waveform abnormalities, and (3) a 2D global model with full-duration, full-lead prototypes for detecting diffuse ECG abnormalities. Each branch was trained independently on a disjoint subset of diagnostic labels, and their outputs were either aggregated via macro-averaging or fused using a learned classifier trained to predict all 71 labels in the PTB-XL dataset.
@@
    Empirical evaluation of prototype quality by clinicians. To assess the interpretability of the learned prototypes, we conducted a structured review in which two physicians independently rated all projected prototypes from the final model on multiple quality criteria, including clarity and class representativeness. This evaluation provides initial evidence that the model’s explanations align with clinical expectations.
@@ 2.2 Contrastive Learning
Our approach is inspired by supervised contrastive learning, which has been widely used to structure encoder representations. SupCon Khosla et al. (2021) and SimCLR Chen et al. (2020) promote similarity among positive pairs and dissimilarity among negatives using log-softmax objectives over large batches. Multi-label extensions, such as MulSupCon Zhang and Wu (2024), weight similarity based on label overlap. However, these instance-level methods are not applicable to our setting, where the goal is to organize a fixed set of prototype vectors with known class assignments—not to embed instances.
@@ 3.1 Dataset and Label Grouping
We used the PTB-XL dataset, a publicly available collection of 21,799 10-second, 12-lead ECGs from 18,869 patients sampled at both 500 Hz and 100 Hz Wagner et al. (2020). Each ECG was annotated with one or more of 71 SCP-ECG labels spanning arrhythmias, conduction disorders, infarction patterns, and morphological abnormalities. We retained the original dataset split from Wagner et al. (2020)—using folds 1-8 for training and fold 9 for validation. We reported performance metrics on fold 10, using it as a hold-out test set. To align with the dataset’s benchmarks, macro-AUROC from Strodthoff et al. (2021) was chosen as the primary performance metric. Appendix D details how ECGs were visualized for figures and prototype review.
The dataset contains several label groupings (e.g., diagnostic superclasses), but these were unsuitable for our desired training process. Consequently, two physicians—one board-certified in cardiology and the other in internal medicine—grouped the 71 labels into three clinically meaningful prototype categories based on the type of visual reasoning required for diagnosis: (1) rhythm-based diagnoses—require temporal pattern analysis across full-length ECG signals, often discernible from a single lead; (2) morphology-based diagnoses—require localized waveform shape or inter-lead comparisons over short time intervals; (3) global diagnoses—require full-lead patterns spanning the full ECG duration. In total, 16 diagnoses were grouped into the 1D rhythm branch, 52 into the 2D morphology branch, and 3 into the 2D global branch. These groupings are detailed in Appendix H.
@@ 3.2 Preprocessing
As the lower end of the frequency range for a normal ECG is 0.5 Hz Zheng et al. (2020), we applied a first-order Butterworth high-pass filter with a 0.5 Hz cutoff. No low-pass filter was applied as we used the 100 Hz samples from the dataset.
@@ 3.3 Model Architecture
This branch was designed to capture long-range temporal dependencies characteristic of rhythm abnormalities. We adopted the ResNet1D-18 architecture from Strodthoff et al. (2021), operating on input ECGs of shape $(12\times 1000)$. The architecture consisted of an initial strided convolution, followed by four residual blocks with increasing channel depth. Feature maps were pooled with an adaptive average pooling layer to produce fixed-length latent representations across leads. These representations formed the prototype matching space for the 1D rhythm model.
@@
To detect localized waveform abnormalities that depend on spatial relationships across ECG leads—such as ST elevation or pathological Q waves—we used a 2D convolutional neural network that treated the 12-lead ECG as a spatial-temporal matrix. Input ECGs were shaped as $(1\times 12\times 1000)$, where the vertical dimension corresponded to the 12 leads and the horizontal dimension to time. The model used a modified ResNet18 He et al. (2015) backbone in which the first convolutional layer was adapted to accept single-channel 2D inputs with a filter size of $(12\times 7)$. Pretrained ImageNet weights were loaded for the remaining layers. To preserve temporal resolution, the global average pooling layer was removed, resulting in a final feature map of shape $(512\times 1\times 32)$, where the temporal axis was downsampled from 1000 to 32 steps. This branch used partial prototypes that spanned a small temporal window (3 units, approximately 0.94 seconds at 100 Hz) and were applied in a sliding fashion across the latent time dimension. During inference, the model computed similarity scores between each partial prototype and all possible time-localized windows of the test ECG’s latent feature map. Top-$k$ pooling (with $k=5$) was then applied to aggregate the $k$ most activated positions into a final similarity score per prototype.
@@
where $\langle\cdot,\cdot\rangle$ denotes the dot product and the scaling factor $a$ adjusts the magnitude of the similarity score based on the latent dimensionality. This cosine similarity variant, adapted from Barnett et al. (2024), was used to compute similarity between each prototype and regions of the latent feature map. For partial prototypes (used in the morphology branch), each prototype slid across the temporal axis of the latent feature map and produced a similarity score at each time step. We then applied top-$k$ pooling (with $k=5$) across these scores to retain the $k$ highest activations, and computed their average to obtain a single similarity score per prototype per input. For global prototypes (used in the rhythm and 2D global branches), each prototype spanned the entire temporal dimension of the latent space, yielding a single similarity score per input without the need for pooling.
@@ 3.4 Prototype Loss Function
Together, our modifications to the clustering and separation loss terms extended their logic from requiring an exact label match between a prototype and a training example to allowing prototypes to be attracted to any example that contains their assigned class—regardless of what other labels are present. This enabled clustering in the presence of multi-label supervision and implicitly supported overlap among frequently co-occurring diagnoses. However, these terms do not directly structure the relationships between prototypes themselves. To more explicitly shape the geometry of the prototype space based on diagnostic co-occurrence, we introduced a contrastive loss applied at the level of prototype-prototype similarity. This encouraged prototypes assigned to frequently co-occurring classes to remain similar in the latent space, while discouraging similarity between prototypes associated with rarely co-occurring diagnoses.
@@
Our contrastive loss was inspired by the Jaccard Similarity Contrastive Loss (JSCL) proposed by Lin et al. (2023) in the context of multi-label text classification. JSCL weights the contrastive objective between sample pairs by the Jaccard similarity of their label sets, allowing for a soft notion of positive and negative pairs in multi-label settings. We adopted this core idea—using Jaccard similarity to scale pairwise contrastive terms—and adapted it to prototype learning by computing similarity between learned prototype vectors. Specifically, we precomputed a Jaccard-based co-occurrence matrix $C\in[0,1]^{P\times P}$ over the training label set and used it to weight the pairwise similarity score between prototypes. $S({p_{i},p_{j}})$ is the similarity between prototype $i$ and prototype $j$, and $P$ is the number of prototypes.
@@ 3.5 Training Procedure
Prototype-based models underwent a three-stage training procedure (see Appendix E). Stage 1 (joint training) and Stage 2 (prototype projection) were repeated as needed (until validation AUC no longer increased) to allow the prototypes to better converge on class-representative features. Stage 3 (fusion classifier training) was conducted only once, after freezing the prototype branches.
3.6 Manual Evaluation of Prototype Quality
To assess the interpretability and perceived clinical utility of the learned prototypes, we conducted a structured review with two practicing physicians. The evaluation focused on the final ProtoECGNet model trained with the full prototype loss suite, including contrastive regularization. Details on the graphical user interface (GUI) used for clinician review and a screenshot of the interface are provided in Appendix C.
Each criterion was rated on a 1–5 Likert scale. Reviewers were instructed to score each prototype independently based on the projected ECG segment (see Appendices C and D for additional details). No test cases or model predictions were presented. ECGs that were identified as having label errors by the clinicians were excluded. We report the mean scores and 95% confidence intervals across reviewers for each evaluation criterion.
4 Results & Discussion
We evaluate ProtoECGNet on classification performance and prototype interpretability. All results are reported on the held-out PTB-XL test set (fold 10). Macro-AUROC is used as the primary evaluation metric throughout to allow comparison to the PTB-XL benchmarking study Strodthoff et al. (2021). However, we also report weighted AUROCs to allow for reporting 95% confidence intervals (CIs) (bootstrapped with 10,000 samples). Weighted AUROCs were required because bootstrapping creates numerous resamplings of the test set, and some of the resamplings do not contain positive examples for all classes (a handful of classes have just one positive example in the test set; sometimes, that one positive example is not included in the resampling). We calculate AUROC for each class in each resampling. For the resamplings that do not contain a positive example of a given class, the AUROC for that class is not defined. The macro-AUROC forms an unweighted average over the AUROCs for each class, and so is also undefined for the resamplings that are missing positive examples of one of the classes. The weighted AUROC averages over the AUROCs in proportion to the number of positive examples in the resampling, and so can safely ignore the AUROCs that are undefined. Class-specific AUROCs for our final model and counts of positive examples for each class are provided in Appendix I.
Experiment: To evaluate whether aligning prototype types with clinical reasoning modalities improves performance, we compare models trained on all 71 PTB-XL labels using (1) a single prototype type (e.g., only 1D or 2D global), and (2) the full ProtoECGNet multi-branch model. Prototype-based models are trained with and without contrastive loss, and the multi-branch outputs are fused using either macro-averaging or a learned fusion classifier.
Results: The best-performing single-branch model was the 2D partial prototype model with contrastive loss, achieving a macro-AUROC of 0.9091 (see Table 1). The full ProtoECGNet model—combining 1D, 2D partial, and 2D global branches—achieved the highest overall performance when using a learned fusion classifier, with a macro-AUROC of 0.9132. This exceeds all single-branch models and approaches the performance of the best models in the PTB-XL benchmarking study (macro-AUROC 0.925 for single model; 0.929 for ensemble) Strodthoff et al. (2021). The same trends were observed when using weighted AUROC (see Table 2); the fusion classifier outperformed single-branch models and achieved 0.9066 (95% CI: 0.9000-0.9128).
Table 1: Macro-AUROC across branch-specific, single-branch, and multi-branch settings for ProtoECGNet.
Setting Model (Label Set) Black-box No Contrastive w/ Contrastive
Branch-Specific Labels Rhythm Branch (16 labels) 0.9403 0.8903 0.9064
Morphology Branch (52 labels) 0.8872 0.8533 0.9051
Global Branch (3 labels) 0.8649 0.8362 0.8667
Full 71-Label, Single Branch 1D Prototype Model 0.9250 0.8646 0.8977
2D Partial Prototype Model N/A 0.8873 0.9091
2D Global Prototype Model 0.8990 0.8681 0.9074
Full 71-Label, Multi-Branch Macro Aggregation 0.8982 0.8609 0.9038
Fusion Classifier N/A 0.8855 0.9132
Table 2: Weighted AUROC with bootstrapped 95% confidence intervals across branch-specific, single-branch, and multi-branch settings.
Setting Model (Label Set) Black-box No Contrastive w/ Contrastive
Branch-Specific Rhythm Branch (16 labels) 0.8919 (0.8757, 0.9071) 0.8762 (0.8616, 0.8901) 0.8853 (0.8708, 0.8991)
Morphology Branch (52 labels) 0.8996 (0.8930, 0.9057) 0.8791 (0.8727, 0.8855) 0.8996 (0.8931, 0.9059)
Global Branch (3 labels) 0.8307 (0.8137, 0.8470) 0.6981 (0.6767, 0.7197) 0.9039 (0.8906, 0.9164)
Single-Branch 1D Prototype Model 0.9081 (0.9012, 0.9147) 0.8108 (0.8025, 0.8188) 0.8857 (0.8782, 0.8930)
2D Partial Prototype Model N/A 0.8605 (0.8526, 0.8684) 0.8743 (0.8666, 0.8819)
2D Global Prototype Model 0.8932 (0.8859, 0.9002) 0.8589 (0.8505, 0.8669) 0.8916 (0.8841, 0.8992)
Multi-Branch Macro Aggregation 0.8855 (0.8779, 0.8926) 0.8486 (0.8413, 0.8557) 0.8950 (0.8882, 0.9016)
Fusion Classifier N/A 0.8800 (0.8728, 0.8872) 0.9066 (0.9000, 0.9128)
Discussion: These results support the hypothesis that modeling distinct diagnostic reasoning types using specialized prototype branches can improve performance. While the contrastive 2D partial prototype model (0.9091 macro-AUROC) was the strongest among single-branch models, the multi-branch ProtoECGNet fusion classifier achieved the highest macro-AUROC (0.9132) and weighted AUROC (0.9066). This demonstrates that structuring the architecture to reflect clinical reasoning modalities does not compromise—and may modestly improve—overall diagnostic performance.
Results: Contrastive loss improved macro-AUROC in all settings. In branch-specific models, the 2D global prototype model improved from 0.8362 to 0.8667, the rhythm model from 0.8903 to 0.9064, and the morphology model from 0.8533 to 0.9051. In the full-label (single-branch) setting, contrastive loss improved the 1D prototype model from 0.8646 to 0.8977, the 2D partial prototype model from 0.8873 to 0.9091, and the 2D global prototype model from 0.8681 to 0.9074. For the multi-branch model, contrastive loss improved macro-AUROC from 0.8855 to 0.9132.
Weighted AUROCs showed similar trends. In the multi-branch fusion model, contrastive loss improved performance from 0.8800 (95% CI: 0.8728-0.8872) to 0.9066 (95% CI: 0.9000-0.9128), with no overlap in confidence intervals. The 2D global branch-specific model also showed a substantial improvement, from 0.6981 (95% CI: 0.6767-0.7197) to 0.9039 (95% CI: 0.8906-0.9164), again with non-overlapping intervals. The morphology branch-specific model increased from 0.8791 (95% CI: 0.8727-0.8855) to 0.8996 (95% CI: 0.8931-0.9059), also without overlap. In contrast, the rhythm branch-specific model improved from 0.8762 (95% CI: 0.8616-0.8901) to 0.8853 (95% CI: 0.8708-0.8991), but with overlapping intervals. The single-branch 1D prototype model improved from 0.8108 (95% CI: 0.8025-0.8188) to 0.8857 (95% CI: 0.8782-0.8930), and the 2D global prototype model from 0.8589 (95% CI: 0.8505-0.8669), both without overlapping intervals. However, confidence intervals slightly overlapped for the single-branch 2D partial prototype model, which improved from 0.8605 (95% CI: 0.8526-0.8684) to 0.8743 (95% CI: 0.8666-0.8819).
Discussion: Contrastive prototype loss improved macro-AUROC across all model settings. Weighted AUROC was also higher for all model configurations, with most settings showing non-overlapping confidence intervals. Notably, contrastive regularization improved both the macro- and weighted AUROC for the final multi-branch ProtoECGNet—supporting the importance of class co-occurrence-informed regularization in multi-label prototype learning.
4.3 How does ProtoECGNet compare to black-box baselines?
Experiment: For each model variant—1D, 2D partial, 2D global, and multi-branch—we compare a black-box ResNet18 baseline with the corresponding prototype models (with and without contrastive loss). This allows us to assess whether our interpretable architecture sacrifices diagnostic performance.
Results: The black-box models achieved macro-AUROCs of 0.9403 (1D rhythm subset), 0.8872 (2D morphology subset), and 0.8649 (2D global subset) when trained on their respective label subsets. When combined, their macro-averaged performance was 0.8982. When trained directly on all 71 labels, the 1D black-box model achieved 0.9250 macro-AUROC and the 2D black-box model achieved 0.8990. Among prototype models trained on all 71 labels, the contrastive 2D partial prototype model achieved 0.9091, and the final contrastive multi-branch fusion model achieved 0.9132.
In terms of weighted AUROC, the 1D black-box model achieved 0.9081 (0.9012, 0.9147), the 2D black-box model 0.8932 (0.8859, 0.9002), and the contrastive fusion model 0.9066 (0.9000, 0.9128). The confidence intervals for the 1D black-box model and the final contrastive fusion model overlapped.
Discussion: The 1D black-box model achieved the highest overall performance (macro-AUROC 0.9250; weighted AUROC 0.9081), matching the strongest non-ensemble model reported in the PTB-XL benchmarking study Strodthoff et al. (2021). The contrastive multi-branch prototype model achieved slightly lower macro-AUROC (0.9132), but its weighted AUROC of 0.9066 had overlapping confidence intervals with the black-box model, indicating statistically comparable performance. These results demonstrate that interpretable, prototype-based models can retain diagnostic accuracy while providing transparent, case-based explanations.
4.4 Are the learned prototypes clinically meaningful?
Experiment: To assess interpretability, we conducted a structured review of all projected prototypes from the final contrastive model. Two physicians—one with board certification in cardiology and one in internal medicine—independently rated each prototype on a 1–5 Likert scale for two criteria: representativeness (how well the prototype exemplifies the assigned class) and clarity (how visually interpretable and artifact-free the segment is). These criteria were adapted from literature on concept-based interpretability (Alvarez-Melis and Jaakkola, 2018; Ghorbani et al., 2019). Specifically, representativeness reflects the “meaningfulness” and “importance” criteria from Ghorbani et al. (2019), which define whether a concept is both human-interpretable and relevant to the model’s prediction. Clarity incorporates aspects of “coherency” from Ghorbani et al. (2019), and “explicitness” from Alvarez-Melis and Jaakkola (2018), capturing how immediately understandable and visually consistent an explanation is.
Reviewer Representativeness (95% CI) Clarity (95% CI)
Cardiologist 4.29 [4.22, 4.35] 4.48 [4.42, 4.54]
Internist 3.59 [3.52, 3.66] 4.73 [4.69, 4.77]
Discussion: These results suggest ProtoECGNet learns meaningful ECG patterns which align with clinician expectations. High average scores for both representativeness and clarity indicate that the prototypes are visually interpretable and diagnostically appropriate. This review provides initial validation that the model generates human-interpretable prototypes while maintaining predictive performance across diverse diagnostic categories but it does not assess downstream utility. Future work plans to include this assessment as a larger user study.
4.5 How does ProtoECGNet support case-based explanations?
4.6 Limitations
One limitation of our work is that our prototype taxonomy relies on a manually defined grouping of PTB-XL labels into rhythm, morphology, and global categories. While this structure was grounded in clinical ECG interpretation heuristics, it introduces simplifications that may not fully capture the nuances of certain diagnoses. Some conditions may exhibit features that span multiple reasoning modalities (e.g., both rhythm and morphology), which are not explicitly modeled in the current architecture. Future work could explore more flexible or data-driven grouping strategies.
Third, we used ResNet-based CNNs as the backbone architecture across all experiments to ensure consistency in comparison. While this design choice was appropriate for isolating the effect of prototype modeling and contrastive loss, it may limit performance relative to more recent architectures. ProtoECGNet is modular by design and compatible with alternative backbones, which should be explored to assess potential gains in performance.
Finally, this study uses only the PTB-XL dataset, which contains ECGs from a single metropolitan area (Berlin, Germany). As such, model generalizability to other populations remains untested. PTB-XL was chosen for its unique diagnostic breadth—71 labels spanning the full clinical scope of ECG interpretation—which enabled development of a method tailored to large-scale, multi-label classification. Nonetheless, external validation on datasets from other settings, even if more limited in scope, will be important in future work to assess broader applicability.
5 Conclusions
Prototype-based learning can be scaled to complex, multi-label medical time-series tasks like ECG classification without sacrificing performance. By separating diagnostic labels into rhythm, morphology, and global categories, ProtoECGNet learned representations aligned with clinically distinct reasoning processes. Each branch learned prototypes tailored to its diagnostic task, and predictions were grounded in real training ECG segments. Our proposed contrastive loss improved performance by structuring the prototype space to reflect label co-occurrence patterns observed in real-world data. Contrastive training consistently outperformed standard prototype objectives and matched black-box baselines. Clinician ratings provided an initial indication that the resulting prototypes were clear and representative of their assigned classes. While our evaluation focused on ECGs, the architectural design and contrastive prototype formulation are broadly applicable to other structured time-series tasks in medicine.
[…]
This work was funded in part by the National Institutes of Health, specifically grant number R00NS114850 to BKB. Additionally, we would like to thank the University of Chicago Center for Research Informatics (CRI) High-Performance Computing team. The CRI is funded by the Biological Sciences Division at the University of Chicago with additional funding provided by the Institute for Translational Medicine, CTSA grant number UL1 TR000430 from the National Institutes of Health.
Appendix A Code Availability
Appendix B ProtoECGNet Internal Architecture
Appendix C Prototype Review Interface
Figure 7: Clinician review interface for prototype evaluation.
Appendix D ECG and Prototype Visualization
Appendix E Training Procedure
Weight Initialization
We initialize the backbones in each branch by first pre-training corresponding blackbox models (i.e. 1D ResNet18, 2D ResNet18) on all 71 labels of PTB-XL. We then initialize a ProtoECGNet, freeze the backbone and classifier layers, and train for 10-20 epochs to “warm up” the prototypes as described in Barnett et al. (2024). This initialization process is optional but can improve performance.